OpenStudy (anonymous):

Sam's plasma TV has a 14 in by 18 in screen with a plastic trim border. The total area of the TV and border is 480 in2. How wide is the trim around the TV?

OpenStudy (anonymous):

480-14*18=the area of the trim=228 let x = trim width. Think of the trim as 8 different parts. One to the right, left, top, and bottom of the screen, and 4 parts in each corner. Each corner is just x^2, and the other four are 14x, 14x, 18x, and 18x. This gives us 228=14x+14x+18x+18x+4*x^2 228=64x+4x^2 57=16x+x^2 Solve using quadratic formula. This is assuming the trim is constant width, and not a function of the height/width.

OpenStudy (anonymous):

I got the same thing, in a very different way. \[(14+x)(18+x)=480\] Here, x is TWICE the x in your equations. That is, it's the total width or height of ALL the trim. So if you solve this, you get x=6, which you have to divide by 2 to get x=3, which solves your equation as well.

OpenStudy (anonymous):

props for subtlety. I like to have every little value, like the area of the screen and stuff. Too much physics :P